OpenStudy (briannabug):

WILL MEADAL AND FAN!!! SOLVE: 3+55x4-5DIVIDED BY 2=?

OpenStudy (briannabug):

@Safa102

OpenStudy (safa102):

You get 213

OpenStudy (welshfella):

recall the order of operations multiply and divide are done before add/subtract

OpenStudy (safa102):

its really easy u just put it in your calculator

OpenStudy (serenity74):

-_-

OpenStudy (safa102):

medal plz........no jk

OpenStudy (safa102):

oh thanks!! :)

OpenStudy (welshfella):

No 213 is incorrect

OpenStudy (safa102):

oh yeah sorry its 108

OpenStudy (safa102):

u had to follow please excuse my dear aunt sally

OpenStudy (briannabug):

ok im confused will u do step by step? please

OpenStudy (welshfella):

3 + 55*4 - 5/2 calculate 55*4 and 5/2 first = 220 and 2.5 so its 3 + 220 - 2.5

OpenStudy (safa102):

u have to follow order of operations first u do multiplication than u divide than u add subtract

OpenStudy (welshfella):

= 223 - 2.5 = 220.5

OpenStudy (briannabug):

yes I got 220.5 thank you for trying @Safa102 thanks for helping me understand @welshfella

OpenStudy (safa102):

np!!

OpenStudy (briannabug):

lol

OpenStudy (welshfella):

yw